OpenStudy (anonymous):

find the slope of the line contaning the points (-1,2) and (5,4)

OpenStudy (anonymous):

1/3.

OpenStudy (aroub):

4-2/5+1 2/6 1/3

OpenStudy (angela210793):

slope=\[(y2-y1)/(x2-x1)\]

OpenStudy (anonymous):

y2-y1/x2-x1 is the formula to calculate slope.

OpenStudy (anonymous):

m=(4-2)/(5-(-1))=2/6=1/3

OpenStudy (anonymous):

GIVE US ALL GOOD ANSWERS! 3:<